DataSpan, Inc., automated its plant at the start of the current year and installed a flexible manufacturing system. The compaRequired: 1-a. Compute the throughput time for each month. 1-b. Compute the delivery cycle time for each month. 1-c. ComputeRequired 1 Required 2 Required 3 1-a. Compute the throughput time for each month. 1-b. Compute the delivery cycle time for eaRequired 1 Required 2 Required 3 Evaluate the companys performance over the last four months. (Indicate the effect of each tRequired 1 Required 2 Required 3 3-a. (Month 5) Refer to the move time, process time, and so forth, given for month 4. Assume
Answer #1

days days Throughput Delivery Cycle Manufacturing Time Time Cycle Efficiency (MCE) Month 1 8.1 days 27. 1 days 28.4 Month 2 2Delivery Cycle time for month 2 = (2018 + 8.6) days e 29.4 days Delivery Cycle time for month 3 = 123.0 + 9.5) days = 32.5 daMonth 5 Mouth 6 Throughput Time 3.3 days 2.9 days - Manufacturing Cycle Efficiency | 60.6 % 69.0 %. Calculation: Bas New Thro